OpenStudy (anonymous):

Identify the factors of 4x2 + 12x + 9. (4x - 3)(x - 3) (4x + 3)(x + 3) (2x - 3)(2x - 3) (2x + 3)(2x + 3)

OpenStudy (anonymous):

Im not sure which one it is

OpenStudy (anonymous):

(4x-3)(x-3) Foil 4x^2-12x-3x-9= 4x^2-9x-9

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

(4x+3)(x+3) Foil 4x^2+12x+3x+9= 4x^2+15x+9

OpenStudy (anonymous):

(2x-3)(2x-3) Foil 4x^2-6x-6x-9= 4x^2-12x-9

OpenStudy (anonymous):

Meaning d?

OpenStudy (anonymous):

the answer is d u are correct

OpenStudy (anonymous):

http://www.hackmath.net/en/calculator/quadratic-equation?a=4&b=12&c=9 so D: 4(x+1.5)(x+1.5) = (2x+3)(2x+3)